Determining the Hardness of Water

(from like a local well)

a) Determine the total hardness of 100ml sample of hard water if 25.6ml of 0.1000M EDTA was titrated into the solution. The solution also contained 2ml of pH 10 ammonium buffer and some eriochrome black indicator for the titration process to occur. Please report this value as mg CaCO3/L. (ppm)

b) Determine the calcium concentration of the same 100ml hard water sample if 20.1ml of 0.1000M EDTA was titrated into the solution. The solution also contained 5ml of diethylamine and some calcon indicator for the titration process to occur.

Multiple titres were done of the same hard water sample to determine total hardness and calcium hardness

c) Determine the magnesium concentration by using the difference between the total hardness and Ca2+.

Numbers and Equations you might need:

CaCO3 + EDTA4- ===> CaEDTA2- + CO32-

EDTA4- + Ca2+ ===> CaEDTA2-

molar mass CaCO3 = 100.09g/mol

molar mass Calcium = 40.08g/mol
Answer #1

Solution:

Total Hardness = 2562 ppm as CaCO3

Calcium Hardness = 2012 ppm as CaCO3, [Calcium concentration = 806 ppm as Ca2+]

Magnesium Hardness = 550 ppm as CaCO3, [Magnesium concentration = 133 ppm as Mg2+]

Details:

From the given equation 1 M of EDTA reacts with 1M of CaCO3 or 1M of Ca2+ or 1M of Mg2+

Therefore for a) 25.6 ml of 0.1M EDTA = 25.6 ml of 0.1M CaCO3 = 25.6 x 0.1 x 100.09 mg as CaCO3

                                                       = 256.2 mg

This much quantity present in 100 ml of the water sample. we know that ppm = mg/L

Therefore 1L of the water sample contains = 256.2 x 1000/100 = 2562 mg or this is equal to 2562 ppm of mg/L

Total Hardness = 2562 ppm as CaCO3

Calcium Hardness =   20.1 x 0.1 x 100.09 x10

                           = 2012 ppm as CaCO3 or = 806 ppm as Ca2+

Magnesium Hardness = [Total Hardness - Calcium Hardness]= 2562-2012 = 0.550 ppm as CaCO3

                                   or = 0.550 x 24.32/100.09 = 0.133 ppm as Mg2+